OpenStudy (anonymous):

What is the value of f(–1) when f(x) = 2x + 2? f(–1) = _________

OpenStudy (anonymous):

\[ f(-1) = 2 \cdot (-1) + 2 = -2 + 2 = \boxed {0}. \]

OpenStudy (anonymous):

all that means is plug -1 for the x values. then solve, very easy

OpenStudy (anonymous):

I know, I posted it, then went back and tried to solve it. I got it right, I have no idea why I just posted it here. :) Thanks anyway!